A manufacturer produces a product which it sells for $60 per unit. The variable cost per unit is $20 and the fixed cost per month is $10,000. 

1.  How many units must be sold per month to break even?

a. 167      b. 500      c. 400      d. 250      e. None of these

 

 

2.  Determine the monthly profit (loss) if it sells 325 units per month.

a. $1500      b. ($1500)      c. $3000      d. $0      e. None of these

 

 

3.      How many units must be sold per month to earn a profit of $7000?

a. 342         b. 425      c. 675          d. 575           e. None of these

 

 

4.  Find the percent discount for an article listed at $40 whose net price is $25.

a. 15%        b. 25%     c. 37.5%      d.  62.5%       e. None of these

 

 

5.   The net price of an article is $125.46. What is the list price if a discount of 38% was allowed?

a. $200.90     b. $204.13       c. $90.91       d. $202.35      e. None of these


1
Expert's answer
2021-02-24T13:50:43-0500

1.

"\frac{1000}{60-20}=250~" $ (d).

2.

"325\cdot(60-20)-10000=3000~" $ (c).

3.

"\frac{7000+10000}{60-20}=425~" $ (b).

4.

"(1-\frac{25}{40})\cdot 100 \%=37.5\%~" (c).

5.

"\frac{125.46}{1-\frac{38}{100}}=202.35~" $ (d).